# Era Splitting - Invariant Learning for Decision Trees

### v4

@@ Abstract
Real-life machine learning problems exhibit distributional shifts in the data from one time to another or from one place to another. This behavior is beyond the scope of the traditional empirical risk minimization paradigm, which assumes i.i.d. distribution of data over time and across locations. The emerging field of out-of-distribution (OOD) generalization addresses this reality with new theory and algorithms which incorporate environmental, or era-wise information into the algorithms. So far, most research has been focused on linear models and/or neural networks. In this research we develop two new splitting criteria for decision trees, which allow us to apply ideas from OOD generalization research to decision tree models, namely, gradient boosting decision trees (GBDT). The new splitting criteria use era-wise information associated with the data to grow tree-based models that are optimal across all disjoint eras in the data, instead of optimal over the entire data set pooled together, which is the default setting. In this paper, two new splitting criteria are defined and analyzed theoretically. Effectiveness is tested on four experiments, ranging from simple, synthetic to complex, real-world applications. In particular we cast the OOD domain-adaptation problem in the context of financial markets, where the new models out-perform state-of-the-art GBDT models on the Numerai data set. The new criteria are incorporated into the Scikit-Learn code base and made freely available online.
@@ 1 Introduction
Gradient boosting regression trees ([30]) and descendent algorithms such as those implemented via the Light Gradient Boosting Machine ([32]), XGBoost ([26]), and Scikit-learn’s HistGradientBoostingRegressor ([38]) are considered state of the art for many real-world supervised learning problems characterized by medium data size, tabular format, and/or low signal-to-noise ratio ([31, 28]). One such data science problem occurs with predicting future returns in the capital markets. When this problem is framed as a standard supervised learning problem, the target variable is assumed to be a continuous score, representing the expected return (alpha) for a particular stock. The input data is the available information about that stock at that particular time. Each stock’s input data and associated target for each date are represented as one row in a data set. This problem is generally referred to as predicting the cross-sectional returns of the equity market.
@@
Supervised learning models in their original form are not aware of the date information presented in the data. One row of input data is associated with one target. Samples are assumed to be drawn I.I.D. from the training and test sets, per the empirical risk minimization (ERM) principal ([40]). While in reality $N$ samples are drawn simultaneously from the data set, where $N$ is the number of stocks in the tradable universe. The predictions for all $N$ stocks are created together at the same time. These $N$ data points may not be completely I.I.D. either, as there are measurable positive correlations among stocks in the stock market over any particular time period. Indeed these ideas are the foundation of the capital asset pricing model (CAPM) ([27]) and modern portfolio theory ([35]), as well as the principles of risk factor models such as the Barra risk model ([24]). Thus, the cross-sectional stock prediction problem violates many of the assumptions of the ERM pricniple that determine the efficacy of the model predictions.
@@
The discrepancy between the foundations of the ERM principle and the realities of ML application fits into the contemporary literature in the academic field of out-of-distribution generalization (OOD) research ([34]). This field of study recognizes that the ERM principle does not hold true for many, if not all, real-life applications. It does not assume that the data distribution of our training and test sets are identical. Instead, the framework assumes a certain distributional shift among disjoint data sets comprising the training data. Models are designed to account for these shifts. In the literature, this concept is often described in terms of environments or domains. In the context of this research, an environment is a domain, is an era, and is a date/location. Drawing data from different environments is analogous to drawing data from different dates or locations. Only data drawn from the same environment follows the same distribution, but across environments the distribution changes. Data sets are composed of data from several environments.
@@
Environments are simply subsets of the entire data set. There is some distributional shift in the data-generating process from one environment to another. The purpose of this is to recognize that there may be spurious signals that work in a single or even a group of environments. It is even shown in ([37]) that these spurious signals can exist when many environments are pooled together. But these spurious signals change or disappear from one environment to another. These spurious signals usually will present themselves as simpler interpretations of the problem, so a naive model will quickly latch onto them. Think of the cow on grass and camel on sand image recognition problem ([36]). During training an image classification task for predicting the label of the animal in an image, cows are always displayed on green grass backgrounds and camels always appear on sandy backgrounds. Naive ML models will simply learn to associate the color green with a cow and the color yellow with a camel. Of course, this is a severe over-simplification of the problem. During evaluation, a cow appearing on a sandy background is classified as a camel, and a camel on a grassy background is classified as a cow. A model which latches onto these spurious signals will perform well in-sample, but have severe flaws when evaluated out-of-sample (OOS).
Fortunately there also exist invariant signals which work in all environments. These signals can often be much more complex than the spurious signals, so naive models are not likely to learn these invariant signals as readily. In the above example, the invariant signal is the cow or camel that is actually present in the image. If an ML model truly understands what a camel or cow looks like, then it will have no problem predicting correctly no matter the background.
@@
The purpose of this field of research is to design ML procedures which allow models to ignore the spurious and learn the invariant signals in data. In the work of ([21]), which is considered foundational in the field, the authors define their problem as a constrained optimization problem and convert it into a loss function that can be used in gradient descent-style training procedures. The loss function combines that tradition ERM risk term, which minimizes the average error over all environments, and an additional penalty on the magnitude of the gradients of the error in each environment. The gradient norm term is used to measure the optimality of the solution in each environment. Considerable theoretical and data analysis is performed around this idea, and the technique is implemented via neural networks and shows that it works on synthetic data problems designed for this purpose, such as the Colored MNIST problem.
@@
A similarly influential work ([37]) also analyzes the phenomenon of spurious and invariant signals existing simultaneously in data, and how to disentangle the signals. Here, the authors look at the loss surface, and describe the co-existence of local minima on the surface where some minima correspond to spurious predictors and others correspond to the invariant predictors. The insight here is to realize that spurious minima occur is some environments and not others, but invariant minima occur in all environments in the same location. The authors extend this reasoning to the directions of the gradients of the loss function with respect to the model parameters. The directions of the gradients for invariant minima should agree across all environments but the directions for spurious signals will disagree between environments. In order to realize this understanding practically, there is an alteration made to the traditional gradient descent algorithm which is called the AND-mask. Before propagating the model weights down the gradients at each iteration, first the gradients are computed using data from each environment separately. Gradient descent is only propagated in the directions of the gradients which agree over all the environments, and not along those that disagree. There is also a variable to adjust the level of agreement required. With this technique the authors show how neural network models can effectively extract invariant predictors from data while ignoring spurious ones. It is important to note that the theoretical insights form the previous two references are limited to the linear settings, while the ML models function in high dimensions and non-linear signals.
The current research article investigates the OOD generalization problem in the context of gradient boosted decision trees (GBDTs), and related decision tree algorithms. In particular we focus on regression trees, which take a continuous variable as target, as opposed to trees for classification which take a class label as target. GBDTs implement decision trees as the base learner $f$. Trees are grown in a top-down approach, which is a cascading, step-wise process which intends to minimize the following objective function at each step, $t\in\{1,2,...\}$ ([25, 39, 30, 26]). The key piece of the objective function is the convex differentiable loss function $l(y_{i},\hat{y_{i}})$ which measures the loss between target $y_{i}$ and prediction $\hat{y}_{i}$. At each step $t$, the loss is computed between the true label and a correction $f_{t}(x_{i})$ to the prediction from the previous step. In the following equations $n$ is the total number of data points for input/target pairs $(x_{i},y_{i}),\forall i\in\{1,...,n\}$ and $\Omega$ is the regularizing function.
$\mathcal{L}^{(t)}=\sum_{i=1}^{n}l(y_{i},\hat{y}^{(t-1)}+f_{t}(x_{i}))+\Omega(f_{t})$ (1)
In practice the second order approximation to this objective is used in optimization, which puts the gradient in gradient boosting. In equation 2 $g_{i}=\partial_{y^{(t-1)}}l(y_{i},\hat{y}^{(t-1)})$ and $h_{i}=\partial^{2}_{y^{(t-1)}}l(y_{i},\hat{y}^{(t-1)})$. This derivation follows closely from ([26]) with more foundation in ([30]) and ([29]).
$\mathcal{L}^{(t)}=\sum_{i=1}^{n}l(y_{i},\hat{y}^{(t-1)})+g_{i}f_{t}(x_{i})+\frac{1}{2}h_{i}f_{t}^{2}(x_{i}))+\Omega(f_{t})$ (2)

### v5

# Era Splitting: Invariant Learning for Decision Trees

@@ Abstract
Real-life machine learning problems exhibit distributional shifts in the data from one time to another or from one place to another. This behavior is beyond the scope of the traditional empirical risk minimization paradigm, which assumes i.i.d. distribution of data over time and across locations. The emerging field of out-of-distribution (OOD) generalization addresses this reality with new theory and algorithms which incorporate environmental, or era-wise information into the algorithms. So far, most research has been focused on linear models and/or neural networks Arjovsky et al. [2020], Parascandolo et al. [2020]. In this research we develop two new splitting criteria for decision trees, which allow us to apply ideas from OOD generalization research to decision tree models, namely, gradient boosting decision trees (GBDTs). The new splitting criteria use era-wise information associated with the data to grow tree-based models that are optimal across all disjoint eras in the data, instead of optimal over the entire data set pooled together, which is the default setting. In this paper, two new splitting criteria are defined and analyzed theoretically. Effectiveness is tested on four experiments, ranging from simple, synthetic to complex, real-world applications. In particular we cast the OOD domain-adaptation problem in the context of financial markets, where the new models out-perform state-of-the-art GBDT models on the Numerai data set. The new criteria are incorporated into the Scikit-Learn code base and made freely available online.
@@ 1 Introduction
Gradient boosting decision trees (GBDTs) Friedman [2001a] and descendent algorithms such as those implemented via the Light Gradient Boosting Machine Ke et al. [2017b], XGBoost Chen and Guestrin [2016a], and Scikit-learn’s HistGradientBoostingRegressor Pedregosa et al. [2011], among others, are considered state of the art for many real-world supervised learning problems characterized by medium data size, tabular format, and/or low signal-to-noise ratio Grinsztajn et al. [2022], Fieberg et al. [2023]. One such data science problem occurs with predicting future returns in the capital markets. When this problem is framed as a standard supervised learning problem, the target variable is assumed to be a continuous score, representing the expected return (alpha) for a particular stock. The input data is the available information about that stock at that particular time. Each stock’s input data and associated target for each date are represented as one row in a data set. This problem is generally referred to as predicting the cross-sectional returns of the equity market.
@@
Supervised learning models in their original form are not aware of the date information presented in the data. One row of input data is associated with one target. Samples are assumed to be drawn I.I.D. from the training and test sets, per the empirical risk minimization (ERM) principal Vapnik [1991]. While in reality $N$ samples are drawn simultaneously from the data set, where $N$ is the number of stocks in the tradable universe. The predictions for all $N$ stocks are created together at the same time. These $N$ data points may not be completely I.I.D. either, as there are measurable positive correlations among stocks in the stock market over any particular time period. Indeed these ideas are the foundation of the capital asset pricing model (CAPM) Fama and French [2004] and modern portfolio theory Markowitz [1952], as well as the principles of risk factor models such as the Barra risk model Blin et al. [2022]. Thus, the cross-sectional stock prediction problem violates many of the assumptions of the ERM principle implicitly assumed in the typical supervised learning setting. Meanwhile, practitioners will often take a GBDT model off the shelf and implement it under these assumptions.
@@
The discrepancy between the foundations of the ERM principle and the realities of ML application fits into the contemporary literature in the academic field of out-of-distribution generalization (OOD) research Liu et al. [2023], Arjovsky et al. [2020], Parascandolo et al. [2020]. This field of study recognizes that the ERM principle does not hold true for many, if not all, real-life applications. It does not assume that the data distribution of our training and test sets are identical. Instead, the framework assumes a certain distributional shift among disjoint data sets comprising the training data. Models are designed to account for these shifts. In the literature, this concept is often described in terms of environments or domains Arjovsky et al. [2020], Peters et al. [2015], Koh et al. [2021]. In the context of this research, an environment is a domain, is an era. Different environments define different experimental conditions that can arise when drawing data at different times or from different locations. Only data drawn from the same environment follows the same distribution, but across environments the distribution can change. Data sets in OOD generalization research are composed of data from several different environments.
@@
There is assumed to be some distributional shift in the data-generating process from one environment to another. The purpose of this is to recognize that there may be spurious signals that work in a single or even a group of environments, but fail to generalize to out-of-sample (OOS) data. It is shown in Parascandolo et al. [2020] that these spurious signals can exist when many environments are pooled together. But these spurious signals change or disappear from one environment to another. These spurious signals usually will present themselves as simpler interpretations of the problem, so a naive model will quickly latch onto them. Think of the cow on grass and camel on sand image recognition problem Nagarajan et al. [2020]. During training an image classification task for predicting the label of the animal in an image, cows are always displayed on green grass backgrounds and camels always appear on sandy backgrounds. Naive ML models will simply learn to associate the color green with a cow and the color of sand with a camel. Of course, this is a severe over-simplification of the problem. During evaluation, a cow appearing on a sandy background is classified as a camel, and a camel on a grassy background is classified as a cow. A model which latches onto these spurious signals will perform well in-sample, but have severe flaws when evaluated OOS. Fortunately there may also exist invariant signals which work in all environments. These signals can often be much more complex than the spurious signals, so naive models are not likely to learn these invariant signals as readily. In this example, the invariant signal is the cow or camel that is actually present in the image. If an ML model truly understands what a camel or cow looks like, then it will have no problem predicting correctly, no matter the background.
Figure 1 shows a conceptual example of what eras represent in financial data. Each era spans a period of time. The boxes spanning each era contain descriptors for the macro-economic environment during each time period. The implication is that financial data coming from each time period (era) contains implicit biases due to the specific conditions of each time period. A question mark is used to show that it is impossible to describe the economic conditions in the future, since that time is still unknown. The traditional ML setting, assuming the ERM principle, implies that the data from the in sample period obey all the same laws as data from the OOS period. OOD research recognizes that this assumption is too idealized for real-life situations, where laws affecting outcomes change from one time to another, or from one place to another.
@@
The purpose of the OOD field of research is to design ML procedures which allow models to ignore the spurious and learn the invariant signals in data, by way of a more accurate representation of the problem. In the work of Arjovsky et al. [2020], which is considered foundational in the field, the authors define their problem as a constrained optimization problem and convert it into a loss function that can be used in gradient descent-style training procedures, such as deep learning with neural networks Goodfellow et al. [2016]. The loss function combines the traditional ERM risk term, which minimizes the average error over all environments, and an additional penalty on the magnitude of the gradients of the error in each environment. The gradient norm term is used to measure the optimality of the solution in each environment. Considerable theoretical and data analysis is performed around this idea, and the technique is implemented via neural networks and shows that it works on synthetic data problems designed for this purpose, such as the Colored MNIST problem.
@@
A similarly influential work Parascandolo et al. [2020] also analyzes the phenomenon of spurious and invariant signals existing simultaneously in data, and how to disentangle the signals with linear models and neural networks. The authors consider the loss surface (the loss function of model parameters), and describe the co-existence of local minima on the surface where some minima correspond to spurious predictors and others correspond to the invariant predictors. The insight here is to realize that spurious minima occur is some environments and not others, but invariant minima occur in all environments in the same location. The authors extend this reasoning and consider the gradients of the loss function with respect to the model parameters. The gradient is constructed as a vector with as many elements as model parameters. The technique is to compute a gradient vector from data in each environment separately. The directions of each gradient element should agree across all environments toward invariant minima but the directions will disagree for spurious signals. In order to realize this understanding practically, the authors alter the traditional gradient descent algorithm by way of the AND-mask Parascandolo et al. [2020]. The AND-mask is a vector the same size as the gradient, containing only ones and zeros (true or false). Elements of the AND-mask corresponding to gradient elements that agree (have the same sign) across all environments are set to 1, and the other elements that disagree (have different signs) are set to zero. The AND-mask is then multiplied element-wise by the gradient of the loss, canceling out elements with disagreement in direction across environments. The masked gradient vector is then used in the downstream gradient descent step, propagating the model parameters toward invariant minima in the loss surface and not toward the spurious minima. The authors then show empirically how neural network models can effectively extract invariant predictors from data while ignoring spurious ones. It is important to note that the theoretical insights form the previous two references are limited to the linear settings, while the neural network models function in high dimensions and non-linear signals.
This current research article investigates the OOD generalization problem in the context of gradient boosted decision trees (GBDTs), and related decision tree algorithms. In particular we focus on regression trees, which take a continuous variable as target, as opposed to trees for classification which take a class label as target. This field of research is largely unexplored, with the only reference to the idea known at the time of writing in Liao et al. [2024], which also looks like a promising direction. It is a related idea but distinct from this paper. The researchers define a criterion to identify invariant splits from the set of potential splits at each node. The predicted values at each node should be consistent over every era of data, to be invariant. The rule is adapted to a regularization term which is combined with the original splitting criterion. The idea is similar, but less explicit, to directional era splitting, which is explained more in detail later. In this research, we do not employ penalties, but instead replace the original split criterion with our own new criteria.
The rest of the introduction offers a review of decision trees for regression and the GBDT algorithm, highlighting the original splitting criterion used in most off-the-shelf GBDT libraries for regression Chen and Guestrin [2016b], Ke et al. [2017a]. The subsequent sections then describe the novel contributions of this research. The era splitting and directional era splitting criteria are presented in sections 2 and 3, which are new contributions of this research. A theoretical breakdown follows in section 4 which offers key insights and motivations to the design choice of this research. Section 5 describes our experimental methods, specifying the OOD details of each of the 4 experiments. Finally, the paper is concluded with a discussion after presenting results. The results indicate that our new splitting criteria lead to better out-of-sample performance and a smaller generalization gap. In particular, directional era splitting stands out as the best. Links to the open source code base are given in the last section of the paper.
1.1 Regression Trees
The basis of the GBDT algorithm for regression are decision trees for regression, also known as regression trees. This section follows closely from the chapter, Regression Trees, of Breiman et al. [1984]. Regression trees are a practical tool that dates back to the 1960s Morgan and Sonquist [1963]. It has been akin to linear regression in spirit and underlying theory, but the implementation details are much different.
The problem setup is common in machine learning and statistics, more generally known as supervised learning. The data consists of pairs of instances, $(x,y)\in(\mathrm{X},\mathrm{Y})$, where $X$ is called the measurement space and $Y$ is the target space. Usually the measurement space is the real numbers of some integer dimension, $d>0$, and the target space is usually just the real numbers, so $(x,y)\in(\mathbb{R}^{d}\times\mathbb{R})$. The $x$ are called the independent (or predictor) variables and the $y$ are the dependent (or response) variables. A prediction rule, or predictor, is a real-valued, parameterized function $f(x,\theta)$ on $\mathbb{R}^{d}$, where $\theta$ represents all the parameters of the predictor. The foundational assumption is made that
$\mathbb{E}[Y|X=x]=f(x,\theta).$ (1)
Regression analysis is the term describing how to construct $f(x,\theta)$ from a training sample $\mathcal{L}$ consisting of $N$ data points $(x_{1},y_{1}),...,(x_{N},y_{N})$. Before getting into the specifics of regression trees as a predictor $f$, let’s first define terms to help us compute the error of a predictor.
Definition 1 (Mean Squared Error, Breiman et al. [1984]).
The mean squared error $R^{*}(f)$ of the predictor $f$ is defined as
$R^{*}(f)=\mathbb{E}[(\mathrm{Y}-f(\mathrm{X}))^{2}],$ (2)
where $(\mathrm{X},\mathrm{Y})$ is an independent sample taken from $\mathcal{L}$.
Given that the training samples are also assumed to be independently sampled from $\mathcal{L}$, the re-substitution estimate, $R(f)$, is the usual method for estimating $R^{*}(f)$, given by
$R(f)=\frac{1}{N}\sum_{n=1}^{N}(y_{n}-f(x_{n}))^{2}.$ (3)
In ML parlance this would be called the training error or loss, since it is the error calculated with the same data that was used to train the model.
The task of regression is to find parameters of a the predictor function of independent variables that best explains a dependent variable in the least squared sense. If $\theta$ is a finite set of parameters, $\theta=(\theta_{1},\theta_{2},...)$, then $\hat{\theta}$ is the parameter value that minimizes the re-substitution error, satisfying
$R(f(x,\hat{\theta}))=\min_{\theta}R(f(x,\theta)).$ (4)
In linear regression, $f$ takes the functional form $f(x,\theta)=\theta_{0}+\theta_{1}*x_{1}+\theta_{2}*x_{2}...$ where $\theta$ is a vector of real-valued numbers to be estimated. In tree regression, $\theta$ represents the tree structure, splitting rules, and parameters that go into defining the tree.
Tree predictors function by splitting the training data into subsets according to some defined splitting rule. The measurement space $\mathrm{X}$ is partitioned by a series of binary splits so that every value of $x\in\mathrm{X}$ falls into a terminal node. Each terminal node is assigned a constant value that acts as the prediction for data instances that end up in that node. In Figure 2, an example tree is drawn, taken from Breiman et al. [1984]. Each node $t_{i}$ for $i\in\{1,2,...,9\}$ is either a parent node or a terminal node. Parent nodes split the data into two child nodes according to the particular split rule for that node. Terminal nodes assign constant values $y(t_{i})$ to the nodes. Any data point falls into one of the terminal nodes of the tree structure, which is assigned that node’s predicted value.
{forest}
[…]
Figure 2: Figure 8.2 Breiman et al. [1984], an example tree.
The three necessary elements for building a regression tree from training data $\mathcal{L}$ are:
  1. 1.
    A method of selecting a split at each node (a splitting criterion)
  2. 2.
    A method to determine if a node is terminal
  3. 3.
    A rule to assign the predicted value $y(t_{i})$ at terminal nodes
An important point to understand is that nodes are said to contain data points. Indeed, trees are grown according to the training data $\mathcal{L}$. The root node, $t_{1}$ always contains all the training data. Split 1 is the rule which partitions all the data into two disjoint subsets, sending the first subset to $t_{2}$ and the second subset to $t_{3}$. Each successive node is then split in a similar fashion, according to the data it contains, until a terminal node is reached. The rest of this section focuses on items 1 and 3 from the list above, defining the splitting criterion of each node and the predicted value of terminal nodes. The method for determining when a node is terminal is something a little less precise. In modern libraries, there are several parameters that control when to stop tree growth, such as strict limits on tree depth or total number of leaves. Growing a tree and then reducing the total number of leaves is also called pruning. A complete treatment of best practices to limit tree growth is beyond the scope of this text. Interested readers are invited to read the reference Breiman et al. [1984] and review the hyper-parameters of XGBoost Chen and Guestrin [2016b] and LightGBM Ke et al. [2017a].
Proposition 1 (Proposition 8.10, Breiman et al. [1984]).
The value of $y(t)$ that minimizes $R(d)$ is the average of $y_{n}$ for all cases $(x_{i},y_{i})$ falling into $t$; that is, the minimizing $y(t)$ is
$\bar{y}(t)=\frac{1}{N(t)}\sum_{x_{i}\in t}y_{i}$ (5)
where the sum is over all $y_{i}$ such that $x_{i}\in t$ and $N(t)$ is the total number of cases in $t$.
Now we take the predicted value in any node $t$ to be $\bar{y}(t)$ and set
$R(T)=\frac{1}{N}\sum_{t\in\tilde{T}}\sum_{x_{i}\in t}(y_{i}-\bar{y}(t))^{2},$ (6)
where $\tilde{T}$ is the set of all terminal nodes, and $\sum_{x_{i}\in t}(y_{i}-\bar{y}(t))^{2}$ is the within node sum of squares. Notice the relation between $R(T)$ and $R(f)$. Given a set $\mathcal{S}$ of potential splits in any node $t\in\tilde{T}$,
Definition 2 (Definition 8.13, Breiman et al. [1984]).
The best split $s^{*}$ of $t$ is that split in $\mathcal{S}$ which most decreases $R(T)$.
The format of this definition that is more common is to consider that any $s\in\mathcal{S}$ of $t$ splits the data into $t_{L}$ (left child node) and $t_{R}$ (right child node), and let
$\Delta R(s,t)=R(t)-R(t_{L})-R(t_{R}),$ (7)
then the best split $s^{*}$ is the one which satisfies
$\Delta R(s^{*},t)=\max_{s\in\mathcal{S}}\Delta R(s,t).$ (8)
The mean-squared error is analogous to the impurity measure for classification trees. The best split is the one which results in the greatest reduction in impurity of the child nodes. In practice, all split rules are defined by one of the independent variables, $x_{i}$, and some real value $v$, such that all data points with $x_{i}<=v$ are sent to $t_{L}$ and the other data points go to $t_{R}$. Define such a split of node $t$ as $s(x_{i},v)$. Thus, the set $\mathcal{S}$ contains all possible splits $s(x_{i},v)$ $\forall$ $i\in N(t)$ and $\forall v\in\mathbb{R}$ and each node. Since data sets are finite, the search for the best $v\in\mathbb{R}$ comes down to searching over the number of distinct values of $x_{i}$ in $\mathcal{L}$.
This subsection described regression trees and how they fit into the field of regression analysis as a type of least squares approximation in function space. We’ve also reviewed how trees are grown, how tree nodes are split, and how predicted values are derived. In the next subsection we will put these regression trees into context with a review GBDTs for regression.
1.2 GBDTs for Regression
The previous section described how to grow a tree on a training data set, by splitting nodes at each step, successively reducing the error of the tree model predictor. GBDTs grow whole trees at each step, resulting in predictors which are a linear combination of many trees. Moreover, each tree is not regressed on the dependent variables $y_{i}$ themselves, but on the gradients of the dependent variables with respect to each data point. This section provides a minimum foundation of GBDTs, following closely from Friedman [2001a], concluding with the definition of the original split criterion Chen and Guestrin [2016b] that is the standard way to score splits in GBDTs for regression.
This section continues with the notation developed in the previous subsection on regression trees. The general machine learning problem, also called the ”predictive learning” problem, is described in Friedman [2001b] as optimal function estimation. The goal is to find an estimate $\hat{F}(x)$ for the function $F^{*}(x)$, which maps from $\mathrm{X}$ to $\mathrm{Y}$, that minimizes the expected value of some loss function over the joint distribution of all $(x,y)\in(\mathrm{X},\mathbb{R})$
$F^{*}=\min_{F}\mathbb{E}_{x,y}R^{*}(F)=\min_{F}\mathbb{E}_{x}\left[\mathbb{E}_{y}\left[R^{*}(F)\right]|x\right].$ (9)
with $R^{*}(F)$ the mean-squared error loss function, as defined in definition 1. GBDTs belong to a class of functions called additive expansions, which are a parameterized class of functions $F(x,\theta)$, with $\theta$ a generic list of parameters, having the form
$F(x,\{\beta_{m},a_{m}\}_{1}^{M})=\sum_{m=1}^{M}\beta_{m}h(x,a_{m}),$ (10)
with $\{\beta_{m},a_{m}\}_{1}^{M}$ the set of parameters and $h(x,a_{m})$ a generic function. In practice $h$ will be a regression tree. This parametric formulation changes the function optimization problem to that of parameter optimization. It is described in Friedman [2001a] how the optimal value of the parameters is often attained by starting with an initial guess, and taking successive steps or boosts based on the sequence of the previous steps toward the optimal value.
Steepest-descent, also known as gradient descent, is one such method. The gradient for element $j$ of the gradient vector at the $m^{\text{th}}$ step is computed as
$\{g_{j,m}\}=\left\{\left[\frac{\partial\mathbb{E}_{x,y}R(F(x,\theta))}{\partial\theta_{j}}\right]_{\theta=\theta_{m-1}}\right\},$ (11)
where $\theta_{m-1}$ are the parameters resulting from the previous step. That parameter is computed as
$\theta_{m-1}=\sum_{i=0}^{m-1}p_{i},$ (12)
where $p_{i}$ are the results of the boosts at each step. These are defined by
$p_{i}=-\rho g_{m}$ (13)
where $g_{m}$ is the vector containing all the $\{g_{j,m}\}$, and $\rho$ is known as the learning rate. This is one technique for numerical optimization in parameter space. An interesting pivot is to consider optimization of a non-parameterized function $F(\mathrm{x})$ in the same way, where the data, $\mathrm{x}$, is taken to be the ”parameter”. The goal is to minimize
$\min_{F}\mathbb{E}_{y}\left[L(y,F(\mathrm{x}))|\mathrm{x}\right]$ (14)
where $L$ is some loss function. Following the numerical optimization with additive expansion paradigm, take the solution to be of the form
$F^{*}(\mathrm{x})=\sum_{m=0}^{M}f_{m}(\mathrm{x}),$ (15)
with $f_{0}(\mathrm{x})$ the initial guess and $f_{i}(\mathrm{x})$ the subsequent boosting rounds for $i\in\{1,2,...,M\}$. For steepest decent we then have
$f_{m}(\mathrm{x})=-\rho g_{m}(\mathrm{x}),$ (16)
with
$g_{m}(\mathrm{x})=\left\{\left[\frac{\partial\mathbb{E}_{y}\left[L(y,F(\mathrm{x}))|\mathrm{x}\right]}{\partial\mathrm{x}}\right]_{F(\mathrm{x})=F(\mathrm{x})_{m-1}}\right\},$ (17)
where
$F_{m-1}=\sum_{i=0}^{m-1}f_{i}(\mathrm{x}).$ (18)
This ideal approach is non-parametric and assumes continuous data. In practice data is finite. With sufficient regularity integration and differentiation can be exchanged, resulting the gradient estimate
$g_{m}(\mathrm{x})=\mathbb{E}_{y}\left[\frac{\partial\left[L(y,F(\mathrm{x}))|\mathrm{x}\right]}{\partial\mathrm{x}}\right]_{F(\mathrm{x})=F(\mathrm{x})_{m-1}}.$ (19)
And so
$F_{m}(\mathrm{x})=F_{m-1}-\rho g_{m-1}(\mathrm{x}).$ (20)
The final step here is a switch back to an assumption of a parameterized additive expansion, as in equation (10). This enforces a certain smoothness over the data, allowing to interpolate values for $x$ lying in between the data points $\{x_{i}\}_{1}^{N}$. The function $h$ from equation (10) is known as the base learner or weak learner. In regression, the base learner is a regression tree, which can define an output in $\mathbb{R}$ for any input in the input space $\mathrm{X}$. The optimization problem becomes a parameter estimation problem again
$\{\beta_{m},a_{m}\}_{1}^{M}=\min_{\hat{\beta}_{m},\hat{a}_{m}}\sum_{1}^{N}L\left(y_{i},\sum_{m=1}^{M}\hat{\beta}_{m}h(x_{i},\hat{a}_{m})\right).$ (21)
The reference Friedman [2001b] then introduces a greedy-stagewise approach to this minimization, for $m\in\{1,2,...,M\}$, with
$(\beta_{m},a_{m})=\min_{\beta,a}\sum_{1}^{N}L\left(y_{i},F_{m-1}(x_{i})+\beta h(x_{i},a)\right)$ (22)
where
$F_{m}(\mathrm{x})=F_{m-1}(\mathrm{x})+\beta h(\mathrm{x},a).$ (23)
This is what is called boosting in the ML literature. Notice the similarity between equations (20) and (23). The function $\beta h(\mathrm{x},a)$ is analogous to the steepest descent step toward $F^{*}(\mathrm{x})$ under the additive expansion assumption of equation (10). However, the the gradients $g_{m}(\mathrm{x})$ can be computed at the data points $x_{i}$, while $h(\mathrm{x},a)$ is a parameterized function that we wish to give a best estimate of $g_{m}(\mathrm{x})$ for all $\mathrm{x}$. In order to do that, the optimal parameters $a_{m}$ for the regression tree $h$ are obtained by
$a_{m}=\operatorname*{arg\,min}_{a,\beta}\sum_{i=0}^{N}[-g_{m}(x_{i})-\beta h(x_{i},a)]^{2}.$ (24)
The gradient boosting model is complete, by defining the update step as follows,
$F_{m}(\mathrm{x})=F_{m-1}(\mathrm{x})+\rho_{m}h(\mathrm{x},a_{m}).$ (25)
In the preceding derivation from Friedman [2001a] leaves out a step from the reference, which optimizes the learning rate(s) $\rho$ for each update. This was left out because in modern ML libraries, the learning rate is usually chosen as a hyper-parameter of the model and kept constant throughout training. Most likely, this preference over time has had to do with the classic bias-variance trade off Hastie et al. [2001], where fitting training data perfectly can be detrimental to model generalization. Usually, optimal parameters are tuned using a hold-out or test set of data or by cross-validation Breiman et al. [1984].
Finally, the first algorithm presented is the gradient boosting algorithm, without tuning the learning rate, in its general form. $L$ is any differentiable loss function, and $\rho$ is the learning rate.
1 $F_{0}(\mathrm{x})=\operatorname*{arg\,min}_{\rho}\sum_{i=1}^{N}L(y_{i},\rho)$
2for $m=1$ to $M$ do
3       for $i=1$ to $N$ do
4             $g_{i}=\left[\frac{\partial\left[L(y_{i},F(x_{i}))|x_{i}\right]}{\partial x_{i}}\right]_{F(\mathrm{x})=F(\mathrm{x})_{m-1}}$
5      $a_{m}=\operatorname*{arg\,min}_{a,\beta}\sum_{i=1}^{N}[g_{i}-\beta h(x_{i},a)]^{2}$
6      $F_{m}(\mathrm{x})=F_{m-1}+\rho h(\mathrm{x},a_{m})$
Algorithm 1 Algorithm 1: Gradient Boost Friedman [2001b]
In the case of least squares regression, when the loss function is $L(y,F(\mathrm{x}))=(y-F(\mathrm{x}))^{2}/2$, then the gradients at each step are just $g_{i}=y_{i}-F_{m-1}(\mathrm{x})$. This finishes our derivation of the gradient boosting algorithm for regression, in the following algorithm.
1 $F_{0}(\mathrm{x})=\operatorname*{arg\,min}_{\rho}\sum_{i=1}^{N}L(y_{i},\rho)$
2for $m=1$ to $M$ do
3       for $i=1$ to $N$ do
4             $g_{i}=y_{i}-F_{m-1}(\mathrm{x})$
5      $a_{m}=\operatorname*{arg\,min}_{a,\beta}\sum_{i=1}^{N}[g_{i}-\beta h(x_{i},a)]^{2}$
6      $F_{m}(\mathrm{x})=F_{m-1}+\rho h(\mathrm{x},a_{m})$
Algorithm 2 Algorithm 2: LS Boost Friedman [2001b]
When the base learner $h$ is a regression tree, then algorithm 2 defines the GBDT for regression algorithm. In this case, notice that line 4 of the algorithm is a least-squares fit of a regression tree via the re-substitution estimate, $R(f)$, of equation (3).
1.3 Splitting